# Learning Communities from Equilibria of Nonlinear Opinion Dynamics

### v1

@@ Abstract
This paper studies community detection for a nonlinear opinion dynamics model from its equilibria. It is assumed that the underlying network is generated from a stochastic block model with two communities, where agents are assigned with community labels and edges are added independently based on these labels. Agents update their opinions following a nonlinear rule that incorporates saturation effects on interactions. It is shown that clustering based on a single equilibrium can detect most community labels (i.e., achieving almost exact recovery), if the two communities differ in size and link probabilities. When the two communities are identical in size and link probabilities, and the inter-community connections are denser than intra-community ones, the algorithm can achieve almost exact recovery under negative influence weights but fails under positive influence weights. Utilizing the fixed point equation and spectral methods, we also propose a detection algorithm based on multiple equilibria, which can detect communities with positive influence weights. Numerical experiments demonstrate the performance of the proposed algorithms.
@@ I-A Related Work
Community detection has been extensively investigated for over two decades [5]. There are three major approaches to the problem. The most common approach involves optimizing quality functions. The Louvain method [11] is a widely-used fast algorithm based on greedy optimization of the modularity, which measure the extent to which a given network partition implies dense connections within subgroups. The Infomap method [12] represents the approach based on dynamics. The method aims to compress trajectories of random walks over networks, by looking for partitions that reduce information required for describing the trajectories. The approach based on statistical inference has become well-established in recent years. The methods infer generative network models that presuppose community structure. The stochastic block model (SBM) is a crucial example, where edges exist with probability depending on pre-assigned community labels. This framework facilitates theoretical analysis of community detectability [13, 14, 15].
@@ II-A Nonlinear Opinion Dynamics
where $d>0$ is the damping coefficient, $u$ is the agent attention parameter to the nonlinear network interaction, and $S$ is an odd saturating function such that $S(0)=0$, $S^{\prime}(0)=1$, and $\textup{sgn}(S^{\prime\prime}(z))=-\textup{sgn}(z)$. Here we assume $S=\tanh$, i.e., the hyperbolic tangent. In the nonlinear term, $\alpha\geq 0$ is the self weight and $\gamma\in\mathbb{R}$ is the influence weight of other agents. Lastly, $b_{i}\in\mathbb{R}$ is an additive input, which can be seen as individual prejudice or external influence.
@@
Suppose that $\mathcal{G}$ is connected, $u\geq 0$, and $\mathbf{b}=0$.
@@
The real-valued opinion $x_{i}$ can be seen as the level of the agent $i$’s support for two options [10, 22, 25]. The sign $\textup{sgn}(x_{i})$ indicates which of the two options the agent supports, and $x_{i}=0$ represents a neutral position. Proposition 1 shows that agreement steady states with all agents having the same sign emerge from the neutral state as $u$ increases, if the influence weight is positive. In contrast, if the influence is negative, disagreement steady states emerge.
@@ IV Detection Algorithms and Main Results
In this section, we first address the community detection problem based on a single equilibrium. The performance of the proposed algorithm is analyzed theoretically for several SBMs (Theorem 1). We then design a detection algorithm based on multiple equilibria, by approximating the adjacency matrix using the fixed point equation of the system and applying spectral clustering techniques to the estimated matrix.
@@ IV-A Detection from Single Equilibrium
Given only one equilibrium, we employ the $k$-means method to cluster the states, as shown in Algorithm 1. Note that the exact equilibrium is assumed to be obtained. We will study detection from noisy observations in the future.
@@
From Proposition 1, the equilibrium $\mathbf{x}^{*}$ reflects the eigenvector centrality if the influence weight is positive. Note that agents in the SSBM have similar centrality. Conditions in (i.a) imply a leader-follower structure, where the community $\mathcal{V}_{2}$ has much fewer agents but denser connections, compared with the community $\mathcal{V}_{1}$. This structure implies different centrality in communities, making detection possible. In the case of negative influence, the equilibrium $\mathbf{x}^{*}$ is close to the eigenvector corresponding to $\lambda_{\min}(A)$. For the disassortative SSBM, the sign of entries of that eigenvector corresponds to community labels, so almost exact recovery can be achieved.
@@ IV-B Detection from Multiple Equilibria
In this subsection, we assume that the system has external inputs, and the resulting equilibria from multiple trajectories are available. Consider the dataset of input-equilibrium pairs $\{[\mathbf{b}^{(1)}~{}\mathbf{x}^{(1)}],\dots,[\mathbf{b}^{(m)}~{}\mathbf{x}^{(m)}]\}$ with $m\in\mathbb{N}_{+}$. We will design a community detection algorithm based on the fixed point equation of the model.
The multiple trajectory case is a commonly studied scenario, in which different episodes of discussion are observed (e.g., [2, 6, 18]). The inputs can be seen as individual prejudice towards different topics or external information that is given for each discussion.

### v2

@@ Abstract
This paper studies community detection for a nonlinear opinion dynamics model from its equilibria. It is assumed that the underlying network is generated from a stochastic block model with two communities, where agents are assigned with community labels and edges are added independently based on these labels. Agents update their opinions following a nonlinear rule that incorporates saturation effects on interactions. It is shown that clustering based on a single equilibrium can detect most community labels (i.e., achieving almost exact recovery), if the two communities differ in size and link probabilities. When the two communities are identical in size and link probabilities, and the inter-community connections are denser than intra-community ones, the algorithm can achieve almost exact recovery under negative influence weights but fails under positive influence weights. Utilizing fixed point equations and spectral methods, we also propose a detection algorithm based on multiple equilibria, which can detect communities with positive influence weights. Numerical experiments demonstrate the performance of the proposed algorithms.
@@ I-A Related Work
Community detection has been extensively investigated for over two decades [5]. There are three major approaches to the problem. The most common approach involves optimizing quality functions. The Louvain method [11] is a widely-used fast algorithm based on greedy optimization of the modularity, which measures the extent a given network partition implying dense connections within subgroups. The Infomap method [12] represents the approach based on dynamics. The method aims to compress random walks over networks, by looking for partitions that reduce information required for describing the trajectories. The approach based on statistical inference has become well-established in recent years. The methods infer generative network models that presuppose community structure. The stochastic block model (SBM) is a crucial example, where edges exist with probability depending on pre-assigned community labels. This framework facilitates theoretical analysis of community detectability [13, 14, 15].
@@ II-A Nonlinear Opinion Dynamics
where $d>0$ is the damping coefficient, $u$ is the agent attention parameter to the nonlinear network interaction, and $S$ is an odd saturating function satisfying $S(0)=0$, $S^{\prime}(0)=1$, and $\textup{sgn}(S^{\prime\prime}(z))=-\textup{sgn}(z)$. Here we assume $S=\tanh$, i.e., the hyperbolic tangent. In the nonlinear term, $\alpha\geq 0$ is the self weight and $\gamma\in\mathbb{R}$ is the influence weight of other agents. Lastly, $b_{i}\in\mathbb{R}$ is an additive input, which can be seen as individual prejudice or external influence.
@@
Suppose that $\mathcal{G}$ is connected, $u\geq 0$, and $\mathbf{b}=\mathbf{0}$.
@@
The real-valued opinion $x_{i}$ represents the agent $i$’s level of support for two options. The sign $\textup{sgn}(x_{i})$ indicates which of the two options the agent supports, and $x_{i}=0$ represents a neutral position. Proposition 1 shows that agreement steady states with all agents having the same sign emerge from the neutral state as $u$ increases, if the influence weight is positive. In contrast, if the influence is negative, disagreement steady states emerge.
@@ IV Detection Algorithms and Main Results
In this section, we first address the community detection problem based on a single equilibrium. The performance of the proposed algorithm is analyzed theoretically for several SBMs (Theorem 1). We then design a detection algorithm based on multiple equilibria, by approximating the adjacency matrix using fixed point equations of the system and applying spectral clustering techniques to the estimated matrix.
@@ IV-A Detection from Single Equilibrium
Given only one equilibrium, we employ the $k$-means method to cluster the states, as shown in Algorithm 1. It is assumed that the exact equilibrium is obtained. We will study detection from noisy observations in the future.
@@
From Proposition 1, the equilibrium $\mathbf{x}^{*}$ reflects the eigenvector centrality if the influence weight is positive. Note that agents in the SSBM have similar centrality. Conditions in (i.a) imply a leader-follower structure, where the community $\mathcal{V}_{2}$ has much fewer agents but denser connections, compared with the community $\mathcal{V}_{1}$. This structure ensures different centrality in communities, making detection possible. In the case of negative influence, the equilibrium $\mathbf{x}^{*}$ is close to the eigenvector corresponding to $\lambda_{\min}(A)$. For the disassortative SSBM, the sign of entries of that eigenvector corresponds to community labels, so almost exact recovery can be achieved.
@@ IV-B Detection from Multiple Equilibria
In this subsection, we assume that the system has external inputs, and the resulting equilibria from multiple trajectories are available. Consider the dataset of input-equilibrium pairs $\{[\mathbf{b}^{(1)}~{}\mathbf{x}^{(1)}],\dots,[\mathbf{b}^{(m)}~{}\mathbf{x}^{(m)}]\}$ with $m\in\mathbb{N}_{+}$. We will design a community detection algorithm based on fixed point equations of the model.
The multiple trajectory case is a commonly studied scenario, in which different discussions are observed (e.g., [2, 6, 18]). The inputs can be seen as individual prejudice towards different topics or external information that is given for each discussion.